# Tunable ultrastrong magnon–magnon coupling and spin Hall magnetoresistance in a van der Waals antiferromagnet

### v3

@@
In Fig. 2(i), we plot the field dependence of ${\Delta}/2\pi$ compared to the half the gaps size as a function of field ($\delta f(H)$) at 15K. We demonstrate a notable difference between the two parameters at the intersection field (0 T when $\theta=45\degree$). Additionally, we observe that ${\Delta}/2\pi$ is field-dependent, as it is determined by the canting angle between the sublattice moments. ${\Delta}/2\pi$ reaches a maximum when the moments are antiparallel at zero field and become zero when the angle exceeds 45 °$\degree$. This contrast with the field-independent ${\Delta}/2\pi$ introduced for the SyAF, where the intrinsic symmetry breaking arises from differing $M_{\mathrm{s}}$ values between the layers [30].
To investigate further, we compare the simulated normalised coupling strength obtained from both methods as a function of $H_{A2}$ (Fig.3(a)). Here, $\eta$ is shown to increase as $H_{A2}$ is decreased, reaching a value of   0.7. At this value, $H_{A2}$ becomes less than $M_{s}$, and the lower mode is coupled strongly enough to be completely suppressed to 0 GHz, at lower values of $H_{A2}$ the gap size can is no longer physical. $\tilde{\Delta}$ increases smoothly, reaching a maximum value of 1.4 when $H_{A2}$ approaches 0 T. Notably, when the lower mode is pushed to 0 GHz, no such discontinuity is observed in $\tilde{\Delta}$. In fact, this corresponds to a value of exactly 1, which can be defined as the critical coupling strength, similar to the deep-strong regime in magnon-photon systems [18, 44]. The critical coupling is defined as the point when the lower hybridised mode is suppressed to 0 GHz at 0 T. Beyond this coupling strength, the lower mode is suppressed to a finite field, where stronger coupling results in a stronger suppressing field.
@@
Eq. 7, provides insight into how to precisely tune the coupling strength through control of the two anisotropy values. In Fig. 3(b), we present a 2D plot of the $\tilde{\Delta}$ for a fixed value of $M_{\mathrm{s}}$ and $H_{\mathrm{E}}$ at 15 K. This demonstrates the possibility to tune the coupling strength from uncoupled to critically coupled regime through manipulation of the two anisotropy terms. Avenues for achieving this precise tuning include ionic gating, intercalation, or strain, all which has been shown to have dramatic effects on anisotropy and magnetisation in vdW magnetic materials [45, 46, 47, 48, 49].
Spin Hall magnetoresistance in a vdW AFM
Figure 3: Spin Hall magnetoresistance in $\mathbf{CrPS}_{4}$ (a) Schematic illustration of SMR in a bilayer device, $J_{c}$ is the applied charge current, $\Delta_{SHE}$ and $\Delta_{iSHE}$ is the change in resistance due to the spin Hall effect (SHE) and inverse spin Hall effect (ISHE), $J_{s}^{abs}$ and $J_{s}^{ref}$ are the absorbed and reflected spin currents respectively. (b) Schematic illustration of the in-plane rotational measurements. The angle $\alpha$ is defined as the angle between the magnetic field (and therefore magnetisation) and the $x$ axis. (c) Atomic force microscopy (AFM) image of the Hall bar device. The inset shows a line scan (red bar) giving the thickness of the full stack $\mathrm{Pt}(3\mathrm{\leavevmode\nobreak\ nm})/\mathrm{CrPS}_{4}(18\mathrm{\leavevmode\nobreak\ nm})/\mathrm{h}-\mathrm{BN}(9.5\mathrm{\leavevmode\nobreak\ nm}$). (d)$R_{\mathrm{xx}}$ signal observed when rotating the sample in-plane under a range of applied fields up to $9\mathrm{\leavevmode\nobreak\ T}$ at 3 K. (e) $\Delta R_{xx}$ (black dots) as a function of applied magnetic field plotted against $M/M_{\mathrm{s}}$ (red). (f-g) $\Delta R_{\mathrm{xx}}$ as a function of temperature. The applied magnetic field is 9 T.
Finally, we present SMR measurements conducted on an exfoliated CrPS4(18 nm) flake on a Pt (3 nm) Hall bar device. The flake is encapsulated with a top layer of h-BN(9.5 nm) to protect it from oxidation effects (Fig. 4(c)). SMR arises in bilayer systems composed of a spin source (Pt) and a magnetic layer (CrPS4) (Fig. 4(a-b)). When a current is passed through the spin source layer in the $x$ direction, a spin current is generated in the $z$ direction with a spin polarization in the $y$ direction. This results in spin accumulation at the interface with the magnetic layer. A transfer of angular momentum will occur with the net magnetic moment if the polarization is non-collinear with the magnetisation, leading to an absorption of the spin current and hence an increase in the total resistance, and vice versa [50, 51]. Thus, by rotating the net moment with an external magnetic field, we observe a longitudinal resistance proportional to $\cos^{2}(\alpha)$. The change in resistance $\Delta R_{xx}$ is given by the following equations:
$R_{xx}=R_{xx0}-\Delta R_{xx}m_{y}{}^{2}$ (8)
where $m_{y}$ is the component of the magnetisation along the $y$ axis. We measure the longitudinal resistance ($R_{xx}$) and transverse resistance ($R_{xy}$) as a function of the angle of the field with respect to the applied current using the 4-point probe method. By comparing the data with a reference sample (Pt), we can confirm whether the measured signal originates from the generated spin current. When comparing the values of the angular-dependent magnetoresistance (ADMR) for out-of-plane rotations (see Fig. S5(a-c), SI), we observe a large magnetoresistance effect in both samples. We attribute this to a significant weak anti-localization (WAL) signal that occurs in Pt at low temperatures and high fields as observed previously by Oyanagi et al.[52] and depicted in more detail in the SI (See Fig. S7). This effect makes it difficult to quantify the SMR signal in this orientation. Instead, we focus on the in-plane orientation and rotate through the angle $\alpha$. In this orientation, it has been shown that we would not expect any signal from WAL [52]. Hanle magnetoresistance could also introduce an SMR-like angular dependence [53]; however, we can eliminate this possibility by comparison with a reference sample of Pt/SiO2. We can confirm in this orientation there is no significant SMR-like signal, and hence we can exclude Hanle magnetoresistance in these devices (see Fig. S7(d), SI).
By measuring the $R_{xx}$ for $\alpha$ rotation at a range of fields (Fig. 4(d)) and temperatures (Fig. 4(f)), we are able to extract the amplitude of the SMR. In Fig. 4(e), the $\Delta R_{xx}$ amplitude is observed to increase linearly at first then saturating at around 8 T, by plotting against the VSM data, we can see that this agrees well with the magnetic moment, indicating that the SMR field dependence in this range is dominated by the magnetic moment as expected. The SMR amplitude observed is the same order of magnitude seen in other vdW and thin film AFM devices [37, 54]. In Fig. 4(g), we plot the temperature dependence of the $\Delta R_{xx}$. We observe a monotonic decrease within the temperature range, approaching zero as we approach $T_{\mathrm{N}}$. Due to the relative thickness of the exfoliated flake, we assume that the $T_{\mathrm{N}}$ in this flake is close to that of the bulk, as we are not within the few-layer regime where the transition temperature will start to drop significantly [55]. The observed temperature dependence is related to the decreasing magnetic moment of the sample as a function of temperature, indicating that this is the dominating factor controlling the SMR amplitude, this is similar to the behaviour observed in other vdW AFM [35]. In addition to this, we observe the presence of a SMR like signal fitting to $\mathrm{cos}(\alpha)\mathrm{sin}(\alpha)$ in $R_{xy}$ (see Fig. S6, SI). This further confirms that Hanle magnetoresistance is not the origin of the SMR signal observed in Fig. 4(d-g). This demonstrates the possibility to read-out the magnetic state of exfoliated flakes through spin current. Furthermore, it allows for future studies using spin-orbit torque (SOT) to switch or manipulate spin textures and magnetisation in exfoliated vdW devices [56].
@@
In conclusion, our investigation has revealed the temperature-dependence of the ultrastrong magnon-magnon coupling in the vdW AFM CrPS4. This coupling is attributed to the interplay of in-plane orthorhombic anisotropy, saturation magnetisation, and the exchange field. We evaluated the normalised coupling strength using two approaches: the gap size $\eta$ and the intrinsic coupling strength $\tilde{\Delta}$. Our findings indicate that, in systems with intrinsic symmetry-breaking anisotropy, these two methods yield notably different normalised coupling strength values. We further provide an analytical description of $\tilde{\Delta}$, showing how tuning anisotropy values beyond a critical point could enhance coupling, similar to the deep-strong coupling regime in magnon-photon systems. Additionally, we demonstrate SMR in an exfoliated Pt/CrPS4 device. Here, we show the feasibility of electrically detecting the magnetic state in such devices, thereby providing a pathway for the electrical readout of semiconducting vdW AFMs. This advancement not only enhances our understanding of the spin dynamics and magneto-transport of vdW AFMs, but also opens avenues for their utilization in next-generation spintronic and magnonic devices.
@@ Acknowledgements
C.W.F.F. and M.C. acknowledge support from the UK National Quantum Technologies programme. C.W.F.F. and H.Y. thank EPSRC for support through EPSRC DTP Case studentship (EP/T517793/1). H.K. acknowledges support from EPSRC via EP/T006749/1 and EP/V035630/1. A.K.B., M.B. and G.E. acknowledges support from the Ministry of Education (MOE), Singapore, under AcRF Tier 3 (MOE2018-T3-1-005). We thank O. Lee for developing Fig. 4(b) and assistance with Blender.
@@
C.W.W.F., H.K., and M.C. conceived and planned the project. C.W.W.F. and M.C. conducted VSM, FMR, and magneto-transport measurements, as well as data analysis. C.W.W.F. conducted AFM measurements. H.Y. and C.W.W.F. carried out MuMax3 simulations in the SI. A.K.B. grew the CrPS4 crystal under the supervision of G.E. and performed XRD measurements under supervision of M.B. Z.X., H.D.L. and T.T. grew platinum films. Z.X. fabricated platinum Hall bars. C.W.F.F. performed transfer of flakes onto devices. C.W.W.F. and M.C. wrote the paper. All authors contributed to discussions on the results and provided feedback on the manuscript.
@@
Correspondence and requests for materials should be addressed to C. W. F. Freeman or M. Cubukcu.

### v4

@@
In Fig. 4(a), we plot the field dependence of ${\Delta}/2\pi$ compared to $g/2\pi$ as a function of field at an example temperature of 15 K. We demonstrate a notable difference between the two parameters at the intersection field (0 T when $\theta=45^{\circ}$). Additionally, we observe that ${\Delta}/2\pi$ is field-dependent, as it is determined by the canting angle between the sublattice moments. ${\Delta}/2\pi$ reaches a maximum when the moments are antiparallel at zero field and becomes zero when the canting angle exceeds $45^{\circ}$. This contrasts with the field-independent ${\Delta}/2\pi$ introduced for the SyAF, where the intrinsic symmetry breaking arises from differing $M_{\mathrm{s}}$ values between the layers and has no field dependence [31].
To further investigate the coupling rate, we compare the simulated coupling rate obtained from both methods as a function of $H_{A2}$, as shown in Fig. 4(b). Here, $\eta$ is observed to increase as $H_{A2}$ decreases, reaching a value of approximately 0.65. At this value, $H_{A2}$ becomes less than $M_{s}$, and micromagnetic simulations show that the system transitions into an in-plane easy-axis system (see SI, Fig. S7). On the other hand, $\tilde{\Delta}$ increases, reaching a maximum value of 1 when $H_{A2}$ equals $M_{\mathbf{s}}$. We define this point as the critical coupling strength, corresponding to the deep-strong regime, similar to observations in anisotropic SyAFs [21, 18, 40].
In Fig. 4(c), we present the micromagnetic simulation of this deep-strong coupling case. In this simulation, the lower hybridized mode is observed to reach 0 GHz at 0 T, indicating the maximum achievable coupling rate. These simulated modes are well fitted by analytical solutions, setting $\mu_{0}H_{A2}$ to $\mu_{0}M_{s}$, which yields a coupling rate of 1.
@@
Eq. 6, provides insight into how to precisely tune the coupling rate through control of the two anisotropy values. In Fig. 4(d), we present a 2D plot of the $\tilde{\Delta}$ as a function of $\mu_{0}H_{A1}$ and $\mu_{0}H_{A2}$ for the values of $\mu_{0}M_{\mathrm{s}}$ and $\mu_{0}H_{\mathrm{E}}$ at 33 K. This demonstrates the possibility to tune the coupling strength from uncoupled to critically coupled regime through manipulation of the two anisotropy terms. Avenues for achieving this precise tuning include ionic gating, intercalation, or strain, all of which have been shown to have considerable effects on anisotropy in vdW magnetic materials [41, 42, 43, 44, 45].
CONCLUSIONS
@@
In conclusion, our investigation has revealed the temperature dependence of the ultra-strong magnon-magnon coupling in the vdW AFM CrPS4, approaching the deep-strong regime with a maximum coupling rate of 0.91 achieved at 33 K. This coupling is attributed to the interplay of orthorhombic anisotropies, saturation magnetisation, and the exchange field. We evaluated the coupling rate using two approaches: the gap size $\eta$ and the intrinsic coupling rate $\tilde{\Delta}$. Our findings indicate that, in systems with intrinsic symmetry-breaking anisotropy, these two methods yield notably different coupling rates. We further provide an analytical description of $\tilde{\Delta}$ in the case of orthorhombic anisotropy, demonstrating how tuning anisotropy values to a critical point could allow for realisation of deep-strong coupling. This study not only enhances our understanding of the spin dynamics in vdW AFMs, but also demonstrates the potential for vdW AFMs as an ideal system to experimentally reach deep-strong magnon-magnon coupling.
@@ ACKNOWLEDGEMENTS
C.W.F.F. and M.C. acknowledge support from the UK National Quantum Technologies and National Measurement System programmes. C.W.F.F. and H.Y. thank EPSRC for support through EPSRC DTP Case studentship (EP/T517793/1). H.K. acknowledges support from EPSRC via EP/T006749/1 and EP/V035630/1. A.K.B., M.B. and G.E. acknowledge support from the Ministry of Education (MOE), Singapore, under AcRF Tier 3 (MOE2018-T3-1-005). We thank O. Lee for assistance with Blender.
@@
C.W.W.F., H.K., and M.C. conceived and planned the project. C.W.W.F. and M.C. conducted VSM, FMR, as well as data analysis. H.Y. and C.W.W.F. carried out MuMax3 simulations. A.K.B. grew the CrPS4 crystal under the supervision of G.E. and performed XRD and SEM with EDS measurements under supervision of M.B.. C.W.W.F. and M.C. wrote the manuscript. All authors contributed to discussions on the results and provided feedback on the manuscript.
@@
Correspondence and requests for materials should be addressed to M.C..